Consider the reaction between sulfuric acid and sodium hydroxide:

H2SO4(aq) + 2 NaOH(aq) → Na2SO4(aq) + 2 H2O(l)                  ΔHrxn = -111 kJ

125 mL of 0.200 M H2SO4(aq) and 125 mL of 0.200 M NaOH(aq) is mixed in a coffee cup calorimeter. Calculate qrxn. The reaction goes to completion.
Answer #1

Answer

qrxn = -1.39 kJ

Explanation

H2SO4(aq) + 2NaOH(aq) --------> Na2SO4(aq) + 2H2O(l) ∆Hrxn = - 111kJ

Stoichiometrically, 1mole of H2SO4 react with 2moles of NaOH to give 2moles of H2O

Given moles of H2SO4 = (0.200mol/1000ml) × 125ml = 0.0250mol

Given moles of NaOH = ( 0.200mol/1000ml) × 125ml = 0.0250mol

0.0250moles of NaOH react with 0.0125moles of H2SO4 to 0.0250moles of H2O

For the formation of 2moles of H2O 111kJ moles of heat is relesaed

Therefore ,

qrxn = (-111kJ/2mol ) × 0.0250mol = -1.39kJ